# Designing A Sustainable Marine Debris Clean-up Framework without Human Labels

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ACM SIGCAS/SIGCHI Conference on Computing and Sustainable Societies; July 08–11, 2024; New Delhi, Indiaprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Computer visionccs: Applied computing
@@ 1. Introduction
The oceans, covering 71% of the earth’s surface, serve as a vast repository of biodiversity and sustain marine life. This precious ecosystem is threatened as various human-made materials, known as ‘marine debris’, enter the ocean. Approximately 75 to 199 tonnes of plastic is currently found in our oceans, resulting in severe environmental, social, economic, and health consequences, according to a report by the United Nations Environment Programme (UNE, 2022). The accumulation of these marine debris objects in a specific region is exacerbated by ocean currents and weather patterns, often linking the marine debris problem to climate change trends for particular oceanic regions (Lincoln et al., 2022). Marine debris includes a spectrum of persistent solid materials such as metals, plastics, rubber, fishing gear, and other discarded items that are manufactured or processed, and abandoned in the marine environment. Government, industry, and community stakeholders are developing marine debris action plans tailored to the local community’s needs, which involve monitoring, coordination, collection, and removal efforts. However, implementing such initiatives can be expensive, both in terms of funding and human efforts.
@@ 2.3. Detecting duplicate objects
Figure 5. Detected clusters of marine debris distribution.
[…]
Detected clusters of marine debris distribution on the island.
@@ 2.4.1. Goal
We used the metadata information from the drone image (longitude, latitude, and altitude) to obtain the real-world point coordinates of the detected objects To map the detected objects to their corresponding location, the spatial resolution can be estimated with the dimension of the bounding box for a detected object. To do this, we needed to calculate the Ground Sample Distance (GSD) for each image location based on the corresponding altitude and camera intrinsic parameters to capture the details.
@@ 3.3. Results
Note that while prior works exist where marine debris objects are detected and classified using computer vision models (Winans et al., 2023; Kikaki et al., 2022; Marin et al., 2021), comparing our results directly with those systems can result in spurious conclusions. First, most prior works trained or finetuned a model using human labels in a supervised learning setting. Second, different marine debris detection involves different numbers and types of object categories depending on the economic activities of the data collection site.

### v2

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: ACM SIGCAS/SIGCHI Conference on Computing and Sustainable Societies; July 8–11, 2024; New Delhi, Indiabooktitle: ACM SIGCAS/SIGCHI Conference on Computing and Sustainable Societies (COMPASS ’24), July 8–11, 2024, New Delhi, Indiadoi: 10.1145/3674829.3675076isbn: 979-8-4007-1048-3/24/07ccs: Computing methodologies Computer visionccs: Applied computing
@@ 1. Introduction
The oceans, covering 71% of the earth’s surface, serve as a vast repository of biodiversity and sustain marine life. This precious ecosystem is threatened as various human-made materials, known as ‘marine debris’, enter the ocean. Approximately 75 to 199 million tons of plastic is currently found in our oceans, resulting in severe environmental, social, economic, and health consequences, according to a report by the United Nations Environment Programme (UNE, 2022). The accumulation of these marine debris objects in a specific region is exacerbated by ocean currents and weather patterns, often linking the marine debris problem to climate change trends for particular oceanic regions (Lincoln et al., 2022). Marine debris includes a spectrum of persistent solid materials such as metals, plastics, rubber, fishing gear, and other discarded items that are manufactured or processed, and abandoned in the marine environment. Government, industry, and community stakeholders are developing marine debris action plans tailored to the local community’s needs, which involve monitoring, coordination, collection, and removal efforts. However, implementing such initiatives can be expensive, both in terms of funding and human efforts.
@@ 2.4.1. Goal
We used the metadata information from the drone image (longitude, latitude, and altitude) to obtain the real-world point coordinates of the detected objects. To map the detected objects to their corresponding location, the spatial resolution can be estimated with the dimension of the bounding box for a detected object. To do this, we needed to calculate the Ground Sample Distance (GSD) for each image location based on the corresponding altitude and camera intrinsic parameters to capture the details.
@@ 2.4.2. Ground Sample Distance
Figure 5. Detected clusters of marine debris distribution.
[…]
Detected clusters of marine debris distribution on the island.
@@ 3.3. Results
Note that while prior works exist where marine debris objects are detected and classified using computer vision models (Winans et al., 2023; Kikaki et al., 2022; Marin et al., 2021), comparing our results directly with those systems can result in spurious conclusions. First, most prior works trained or finetuned a model using human labels in a supervised learning setting. Second, different marine debris detection methods involves different numbers and types of object categories depending on the economic activities of the data collection site.